What is the sum of the measures of the interior angle of a polygon with 7 sides?

Answer by ewatrrr(24785)

 
Hi
What is the sum of the measures of the interior angle of a polygon with 7 sides?
Note: In general 180°(n-2) is the sum of the interior angles of a polygon with n-sides
180°(7-2) = 180*5 = 900°